OpenStudy (anonymous):

Please help me? :) My answer: 7 - 3i over 58 Sq root -1 over (2-4i) - (5+3i) =?

OpenStudy (anonymous):

and if your question says \[\frac{ \sqrt{-1} }{ -3-7i }=\frac{ -i }{3+7i }*\frac{ 3-7i }{ 3-7i }=\frac{ 7-3i}{ 9+49 }\]

OpenStudy (anonymous):

your right